OpenStudy (anonymous):

expand sinx in powers of (x- pi/2) and then using it to approximate sin 91 degree!! Help plzzz!!

OpenStudy (anonymous):

i already expanded the series! but cannot understand how to find the value of sin 91 degree!!

OpenStudy (anonymous):

\[sinx = 1-\frac{ (x-\frac{ \Pi }{ 2 })^{2}}{2!}+\frac{ (x-\frac{ \Pi }{ 2 })^{4}}{4!}-......\]

OpenStudy (anonymous):

@wio help plzzz!!

OpenStudy (anonymous):

Um, \(x=91\)

OpenStudy (anonymous):

but it is in degrees!! how to put the value??

OpenStudy (anonymous):

\[ 91^\circ = \frac{91\pi}{180} \]

OpenStudy (anonymous):

yeah then what would be the value?? would it be 1.588 ??

OpenStudy (anonymous):

Just plug it in.

OpenStudy (anonymous):

thanks man!! appreciate it!!